A spaceman jumps from an airplane. After he had fallen 35 m, then his parachute opens. Now he falls with retardation of 1.8 m/s² and reaches the earth with a velocity of 2.7 m/s. What was the height of the airplane?
1
Expert's answer
2020-05-03T17:17:13-0400

The speed of a spaceman after he had fallen 35 m

"v_i=\sqrt{2gh}=\sqrt{2\times 9.8\times 35}=26.2\:\rm m/s"

The distance traveled by spaceman when he falls with retardation

"d=\frac{v_f^2-v_i^2}{2a}=\frac{2.7^2-26.2^2}{2\times (-1.8)}=189\:\rm m"

Finally, the height of the airplane

"H=h+d=35+189=224\:\rm m."